OpenStudy (anonymous):

A circle has a diameter of 10 millimeters. What are the circumference and area of the circle? Use for 3.14 for pi . Round to the nearest tenth.

OpenStudy (lgbasallote):

Circumference = \(\pi d\) = 3.14 x diameter Area = \(\pi (d/2)^2\) = 3.14 x (d/2)^2 got it?

OpenStudy (anonymous):

No

OpenStudy (lgbasallote):

it's the formula for circumference and area

OpenStudy (anonymous):